# Customizing Text-to-Image Diffusion with Camera Viewpoint Control

### v1

@@ Abstract
Model customization introduces new concepts to existing text-to-image models, enabling the generation of the new concept in novel contexts. However, such methods lack accurate camera view control w.r.t the object, and users must resort to prompt engineering (e.g., adding “top-view”) to achieve coarse view control. In this work, we introduce a new task – enabling explicit control of camera viewpoint for model customization. This allows us to modify object properties amongst various background scenes via text prompts, all while incorporating the target camera pose as additional control. This new task presents significant challenges in merging a 3D representation from the multi-view images of the new concept with a general, 2D text-to-image model. To bridge this gap, we propose to condition the 2D diffusion process on rendered, view-dependent features of the new object. During training, we jointly adapt the 2D diffusion modules and 3D feature predictions to reconstruct the object’s appearance and geometry while reducing overfitting to the input multi-view images. Our method outperforms existing image editing and model personalization baselines in preserving the custom object’s identity while following the input text prompt and the object’s camera pose.
Figure 1: Given multi-view images of a new object (left), we create a customized text-to-image diffusion model with camera pose control. The customized model can synthesize the object (denoted as V ¡category name¿) in novel appearances and scenes, such as A green V car, or A beetle-like V car, while allowing users to specify the target camera pose for the object. We can generate panorama images while controlling the camera pose for each object or compose multiple concepts using MultiDiffusion [3] with our model.
@@ 1 Introduction
Recently, we have witnessed an explosion of works on customizing text-to-image models [74, 23, 45, 16]. They allow us to quickly acquire visual concepts, such as personal objects and favorite places, and reimagine them with new environments and attributes. For instance, we can customize a model on our Teddy bear and prompt it with “Teddy bear on a bench in the park.” Unfortunately, customization methods lack precise camera pose control, as existing diffusion models are trained purely on 2D images without ground truth camera poses. As a result, users often rely on text prompts such as “front-facing” or “side-facing”, a tedious and unwieldy process to control views.
@@
What if we wish to synthesize a new object, e.g., the Teddy bear in Figure 1, in a different context while controlling its pose? In this work, we introduce a new task: given multi-view images of the object, we customize a text-to-image model while enabling precise control of the new object’s camera pose. During inference, our method offers the flexibility of conditioning the generation process on both a target pose and a text prompt.
@@
Neural rendering methods have allowed us to accurately control the 3D viewpoint of an existing scene, given multi-view images [39, 57, 5, 4]. Similarly, we seek to imagine the object from novel viewpoints but in a new scene. However, as pre-trained diffusion models, such as Latent Diffusion models [72], are built upon a purely 2D representation, connecting the 3D neural representation of the object to the 2D internal features of the diffusion model remains challenging.
@@
In this work, we propose a new method, CustomDiffusion360, to bridge the gap between 3D neural capture and 2D text-to-image diffusion models by providing additional camera pose control w.r.t. the new custom object in 2D text-to-image models. More concretely, given multi-view images of an object, we learn to predict neural feature fields in the intermediate feature spaces of the diffusion model U-Net. To condition the generation process on a target pose, we render the features of the predicted feature fields and then fuse them with the target pose’s noisy features. This new conditioning module is added to a subset of transformer layers in the pre-trained diffusion model. We only train the parameters of the new feature prediction module to preserve object identity and increase generalization. All parameters of the pre-trained model remain frozen, thus keeping our method computationally and storage efficient.
@@
We build our method on Stable Diffusion-XL [64] and show results on various object categories, such as cars, chairs, motorcycles, teddy bears, and toys. We compare our method with image editing, model customization, and NeRF editing methods. Our method maintains high alignment with the target object and poses while adhering to the user-provided text prompt. We show that directly integrating the 3D object information into the text-to-image model pipeline provides performance gains relative to existing 2D methods. Our method can also be combined with other algorithms [52, 3] for applications like generating objects in different target camera poses while preserving the background, panorama synthesis, or composing multiple concepts. Our code, results, and data are available on our webpage.
@@ 2 Related Works
Text-based image synthesis.   Large-scale text-to-image models [77, 69, 22, 34, 99] have become ubiquitous with their capabilities of generating photorealistic images from text prompts. This progress has been driven by the availability of large-scale datasets [80] as well as advancements in model architecture and training objectives [19, 79, 35, 63, 36]. Among them, diffusion models [85, 30, 83] have emerged as a powerful family of models that generate images by gradually denoising Gaussian noise. Their learned priors have been found useful in various applications, such as image editing [52, 29, 102] and 3D creation [65, 47].
@@
Image editing.   One crucial application enabled by the above models is image editing based on text instructions [61]. For example, SDEdit [52] exploits the denoising nature of diffusion models, guiding generation in later denoising timesteps using edit instructions while preserving the input image layout. Various works aim to improve upon this by embedding the input image into the model’s latent space [85, 38, 55, 60], while some use cross-attention and self-attention mechanism for realistic and targeted edits [29, 13, 25, 62, 10]. Recently, several methods train conditional diffusion models to follow user edit instructions or spatial controls [102, 9]. However, existing methods primarily focus on changing style and appearance, while our work enables both viewpoint and appearance control.
@@
Model customization.  While pre-trained models are trained to generate common objects, users often wish to synthesize images with concepts from their own lives. This has given rise to the emerging technique of model personalization or customization [74, 23, 45]. These methods aim at embedding a new concept, e.g., pet dog, toy, personal car, person, etc., into the output space of text-to-image models. This enables generating new images of the concept in unseen scenarios using the text prompt, e.g., my car in a field of sunflowers. To achieve this, various works fine-tune a small subset of model parameters [45, 26, 32, 89] and/or optimize text token embeddings [23, 92, 104, 1] on the few images of the new concept with different regularizations [74, 45]. More recently, several encoder-based methods have been proposed that train a model on a vast dataset of concept library [81, 2, 24, 94, 46, 90, 75], enabling faster customization during inference. However, none of the existing works allow controlling the camera pose during inference time. In contrast, given the ease of capturing multi-view images of a new concept, in this work, we ask whether we can augment the capabilities of model customization with additional control of the camera pose.
@@
Figure 2: Overview. We propose a model customization method that utilizes $N$ reference images defining the 3D structure of an object $\mathcal{Y}$ (we illustrate with 2 views for simplicity). We modify the diffusion model U-Net with pose-conditioned transformer blocks. Our Pose-conditioned transformer block features a FeatureNeRF module, which aggregates features from the individual viewpoints to target viewpoint $\phi$, as shown in detail in Figure 3. The rendered feature $W_{y}$ is concatenated with the target noisy feature $W_{\mathbf{x}}$ and projected to the original channel dimension. We use the diffusion U-Net itself to extract features of reference images, as shown in the top row. We only fine-tune the new parameters in linear projection layer $l$ and FeatureNerF in $F_{\text{pose}}$ blocks.
@@
View synthesis.   Novel view synthesis aims to render a scene from unseen camera poses, given multi-view images. Recently, the success of volumetric rendering-based approaches like NeRF [54] have led to numerous follow-up works with better quality [4, 5], faster speed [57, 14], and fewer training views [98, 58, 18, 86]. Recent works learn generative models with large-scale multi-view data to learn generalizable representations for novel view synthesis [49, 78, 106, 11, 95, 50]. While our work draws motivation from this line of research, our goal differs - we aim to enable 3D control in text-to-image personalization, rather than capturing real scenes. Recently, Cheng et al. [17] and Höllein et al. [31] propose adding camera pose in text-to-image diffusion models, while we focus on model customization.
@@
3D editing.   Loosely related to our work, many works have been proposed for inserting and manipulating 3D objects within 2D real photographs, using classic geometry-based approaches [37, 15, 41] or recent generative modeling techniques [96, 103, 53]. Instead of editing a single image, our work aims to “edit” the model weights of a pre-trained 2D diffusion model. Many recent works edit [27, 20] or generate [68, 88, 82] a 3D scene given a text prompt or an image. These methods focus on ensuring the multi-view consistency of the scene. Unlike these, we do not aim to edit a 3D multi-view consistent scene, but instead provide additional camera pose control for the new object when customizing text-to-image models.
@@ 3 Method
Given multi-view images of an object, we aim to embed the object in the text-to-image diffusion model. We construct our method in order to allow the generation of new variations of the object through text prompts while providing control of the camera pose of the object. Our approach involves fine-tuning a pre-trained text-to-image diffusion model while conditioning it on a 3D representation of the object learned in the diffusion model’s feature space. In this section, we briefly overview the diffusion model and then explain our method in detail.
@@ 3.1 Diffusion Models
where $\mathbf{c}$ can be any modality to condition the generation process. The model is trained to predict the noise added to create the input noisy image ${\mathbf{x}}_{t}$. During inference, we gradually denoise a random Gaussian noise for a fixed set of timesteps. Various sampling strategies [85, 51, 35] have been proposed to reduce the number of sampling steps $T$ compared to the usual $1000$ timesteps in training.
3.2 Customization with Camera Pose Control
@@
Model customization aims to condition the model on a new concept, given $N$ representative images of the concept $\mathcal{Y}=\{{\mathbf{y}}_{i}\}_{i=1}^{N}$, i.e., to model $p({\mathbf{x}}|\mathcal{Y},\mathbf{c})$ with text prompt $\mathbf{c}$. In contrast, we aim to additionally condition the model on camera pose, allowing more control in the generation process. Thus, given a set of multi-view images $\{{\mathbf{y}}_{i}\}_{i=1}^{N}$ and the corresponding camera poses $\{\pi_{i}\}_{i=1}^{N}$, we seek to achieve a customized text-to-image model corresponding to the object, i.e., our goal is to learn the conditional distribution $p({\mathbf{x}}|\{({\mathbf{y}}_{i},\pi_{i})\}_{i=1}^{N},\mathbf{c},\phi)$, where $\mathbf{c}$ is text prompt and $\phi$ is the target camera pose. To achieve this, we fine-tune a pre-trained text-to-image diffusion model, which models $p({\mathbf{x}}|\mathbf{c})$, with the additional conditioning of target pose and reference views.
@@
Model architecture.   In Figure 2, we show our architecture, with an emphasis on the added pose-conditioned transformer block. To begin, we use the Stable Diffusion-XL (SDXL) [64] as the pre-trained text-to-image diffusion model in our work. It is based on the Latent Diffusion Model (LDM) [72], which is trained in an autoencoder [43] latent space. The diffusion model is a U-Net [73] consisting of encoder, middle, and decoder blocks. Each block consists of a ResNet [28], denoted as $h$, followed by several transformer layers [91]. Each transformer layer consists of a self-attention layer (denoted as $s$), followed by a cross-attention (denoted as $g$) with the text condition, and a feed-forward MLP (denoted as $f$). Given feature map ${\mathbf{z}}$, output of an intermediate ResNet layer $h$ in the U-Net, a standard transformer block performs $F_{\text{standard}}({\mathbf{z}},\mathbf{c})=f(g(s({\mathbf{z}}),\mathbf{c}))$. We modify the transformer layer to incorporate pose conditioning.
@@
Pose-conditioned transformer block.   To condition the model on the 3D structure of the object, we modify the transformer block to be a pose-conditioned transformer block $F_{\text{pose}}({\mathbf{z}}_{0},\{{\mathbf{z}}_{i},\pi_{i}\},\mathbf{c},\phi)$, where ${\mathbf{z}}_{0}$ is the feature from the main branch, $\{{\mathbf{z}}_{i}\}$ are intermediate feature maps corresponding to multi-view images, and $\{\pi_{i}\}$ and $\phi$ are reference and target camera poses. To condition on the multi-view images, we learn a radiance field conditioned on reference view features in a feed-forward manner [98]. We extract features $\{{\mathbf{W}}_{i}\in\mathbb{R}^{H\times W\times C}\}$ from $\{{\mathbf{z}}_{i}\}$. We use components of pre-trained U-Net itself, $F_{\text{standard}}$, to extract these features and render them into a target pose $\phi$ using a FeatureNeRF function to obtain 2D feature map ${\mathbf{W}}_{y}$.
$\begin{split}{\mathbf{W}}_{i}=F_{\text{standard}}({\mathbf{z}}_{i},\mathbf{c}),\hskip 5.69054pt{\mathbf{W}}_{y}=\text{FeatureNeRF}(\{{\mathbf{W}}_{i},\pi_{i}\},\mathbf{c},\phi)\end{split}$ (3)
FeatureNeRF.   Here, we describe the aggregation of individual 2D features ${\mathbf{W}}_{i}$ with 3D poses $\pi_{i}$ into a feature map ${\mathbf{W}}_{y}$ from pose $\phi$. Rather than learning NeRF in a feature space [97, 40], our focus is on learning 3D features that the 2D diffusion model can use. From target viewpoint $\phi$, for each point $\mathbf{p}$ on a target ray with direction $\mathbf{d}$, we project the point on the image plane of given views $\pi_{i}$ and denote projected locations as $\pi_{i}^{\mathbf{p}}$. We then sample from these coordinates on feature map ${\mathbf{W}}_{i}$, predict a feature for the 3D point, and aggregate them with function $\psi$:
$\mathbf{V}_{i}=$ $\text{MLP}(\text{Sample}({\mathbf{W}}_{i};\pi_{i}^{\mathbf{p}}),\gamma(\mathbf{d}),\gamma(\mathbf{p})),\;{i=1,...,N}$ (4)
$\bar{\mathbf{V}}=\psi(\mathbf{V}_{1},...,\mathbf{V}_{N}),$
@@
Figure 3: FeatureNeRF block. We predict volumetric features $\bar{\mathbf{V}}$ for each 3D point in the grid using reference features $\{W_{i}\}$ (Eqn. 4). Given this feature, we predict the density $\sigma$ and color $rgb$ using a 2-layer MLP and use the predicted density to render $\hat{\mathbf{V}}$, which has been updated with text cross-attention $g$. The predicted $rgb$ is only used to calculate reconstruction loss during training.
@@
where $\gamma$ is the frequency encoding. We use the weighted average [71] as the aggregation function $\psi$, where a linear layer predicts the weights based on $\mathbf{V}_{i}$, $\pi_{i}$, and target pose $\phi$. For each reference view, $\mathbf{d}$ and $\mathbf{p}$ are first transformed in the view coordinate space. We then predict the density and color of the point using a linear layer:
$(\sigma,\mathbf{C})=\text{MLP}(\bar{\mathbf{V}}).$ (5)
@@
We also update the aggregated feature with text condition $\mathbf{c}$ using cross-attention:
$\hat{\mathbf{V}}=\text{CrossAttn}(\bar{\mathbf{V}},\mathbf{c}),$ (6)
@@
and render this updated intermediate feature using the predicted density $\sigma$:
@@
where $r$ is the target ray, $\hat{\mathbf{V}}$ is the aggregated feature of the point in the ray, $\sigma$ is the predicted density of that point, $N_{f}$ is the number of sampled points along the ray between the near and far plane of the camera, and $T_{j}=\exp(-\sum_{k=1}^{j-1}\sigma_{k}\delta_{k})$ handles occlusion until that point.
Conditioning.   To process the main target branch, we extract the intermediate 2D feature map after the self $s$ and cross-attention layers $g$, i.e., ${\mathbf{W}}_{{\mathbf{x}}}=g(s({\mathbf{z}}_{0}),\mathbf{c})$. We concatenate ${\mathbf{W}}_{{\mathbf{x}}}$ with the rendered features ${\mathbf{W}}_{y}$ and then project back into the original feature dimension using a linear layer. Thus, the output of the modified transformer layer is:
$F_{\text{pose}}=f(l({\mathbf{W}}_{y}\oplus{\mathbf{W}}_{{\mathbf{x}}})),$ (8)
where $l$ is a learnable weight matrix, which projects the feature into the space to be processed by the feedforward layer $f$. We initialize $l$ such that the contribution from ${\mathbf{W}}_{y}$ is zero at the start of training.
@@
and two mask-based losses – (1) silhouette loss [70] $\mathcal{L}_{\text{s}}$ which forces the rendered opacity to be similar to object mask, and (2) background suppression loss [6, 7] $\mathcal{L}_{\text{bg}}$ which enforces the density of all background rays to be zero, as we only wish to model the object.

### v2

# Customizing Text-to-Image Diffusion with Object Viewpoint Control

@@ Abstract
Model customization introduces new concepts to existing text-to-image models, enabling the generation of these new concepts/objects in novel contexts. However, such methods lack accurate camera view control with respect to the new object, and users must resort to prompt engineering (e.g., adding “top-view”) to achieve coarse view control. In this work, we introduce a new task – enabling explicit control of the object viewpoint in the customization of text-to-image diffusion models. This allows us to modify the custom object’s properties and generate it in various background scenes via text prompts, all while incorporating the object viewpoint as an additional control. This new task presents significant challenges, as one must harmoniously merge a 3D representation from the multi-view images with the 2D pre-trained model. To bridge this gap, we propose to condition the diffusion process on the 3D object features rendered from the target viewpoint. During training, we fine-tune the 3D feature prediction modules to reconstruct the object’s appearance and geometry, while reducing overfitting to the input multi-view images. Our method outperforms existing image editing and model customization baselines in preserving the custom object’s identity while following the target object viewpoint and the text prompt.
Figure 1: Given multi-view images of a new object (left), denoted as V ¡category name¿, we create a customized text-to-image diffusion model with object viewpoint control. The customized model allows users to specify the target viewpoint for the object while synthesizing it in novel appearances and scenes, such as A green V car, or A beetle-like V car. We can also generate panorama images or compose multiple concepts while controlling each object’s viewpoint by using MultiDiffusion [4] with our model.
@@ 1 Introduction
Recently, we have witnessed an explosion of works on customizing text-to-image models [74, 25, 47, 18]. Such methods enable a model to quickly acquire visual concepts, such as personal objects and favorite places, and re-imagine them with new environments and attributes. For instance, we can customize a model on our teddy bear and prompt it with “Teddy bear on a bench in the park.” However, these methods lack precise viewpoint control, as the pre-trained model is trained purely on 2D images without ground truth camera poses. As a result, users often rely on text prompts such as “front-facing” or “side-facing”, a tedious and unwieldy process to control views.
@@
What if a user wishes to control the custom object’s viewpoint while synthesizing it in a different context, e.g., the car in Figure 1? In this work, we introduce a new task: given multi-view images of an object, we customize a text-to-image model while enabling control of the object’s viewpoint. During inference, our method offers the flexibility of conditioning the generation process on both a target viewpoint and a text prompt.
@@
Neural rendering methods have allowed us to accurately control the 3D viewpoint of an existing scene, given multi-view images [41, 59, 6, 5]. Similarly, we seek to imagine the object from novel views but in a new context. However, as pre-trained diffusion models, such as Latent Diffusion models [72], are built upon a purely 2D representation, connecting the 3D neural representation of the object to the 2D internal features of the diffusion model remains challenging.
@@
In this work, we introduce CustomDiffusion360, a new method to bridge the gap between 3D neural capture and 2D text-to-image diffusion models by providing viewpoint control for custom objects. More concretely, given multi-view images of an object, we introduce FeatureNeRF blocks in the diffusion model U-Net’s intermediate feature spaces to learn view-dependent features. To condition the generation process on a target viewpoint, we render the FeatureNeRF output from this viewpoint and merge it with the diffusion features using linear projection layers. We only train the new linear projection layers and FeatureNeRF blocks, added to a subset of transformer layers, to preserve object identity while maintaining generalization. The pre-trained model’s parameters remain frozen, thus keeping our method computationally and storage efficient.
@@
We build our method on Stable Diffusion-XL [65] and show results on various object categories, such as cars, chairs, motorcycles, teddy bears, and toys. We compare our approach with image editing [53, 10], model customization [34], and NeRF editing methods [22]. Our method achieves high alignment with the custom object’s identity and target viewpoint while adhering to the user-provided text prompt. We show that integrating the 3D object information into the text-to-image model, as done by our method, enhances performance over 2D and 3D editing baseline methods. Additionally, our method can be combined with other algorithms [53, 4] for applications such as object viewpoint adjustment in the same background, panorama synthesis, and object composition.
@@ 2 Related Works
Text-based image synthesis.   Large-scale text-to-image models [77, 69, 24, 36, 103] have become ubiquitous for generating photorealistic images from text prompts. This progress has been driven by the availability of large-scale datasets [80] as well as advancements in model architecture and training objectives [21, 79, 37, 64, 38]. Among them, diffusion models [85, 32] have emerged as a powerful family of models that generate images by gradually denoising Gaussian noise.
@@
Image editing with text-to-image diffusion.   One of the first works, SDEdit [53], exploited the denoising nature of diffusion models, guiding generation in later denoising timesteps using edit instructions while preserving the input image layout. Since then, various works improved upon this by embedding the input image into the model’s latent space [85, 40, 57, 62] or using cross-attention and self-attention mechanisms for realistic and targeted edits [31, 15, 27, 63, 12]. Recently, several methods train conditional diffusion models to follow user edit instructions or spatial controls [107, 10]. However, these methods primarily focus on appearance editing, while our work enables both viewpoint and appearance control.
@@
Model customization.  While pre-trained models can generate common objects, users often wish to synthesize images with concepts from their own lives. This has given rise to the emerging technique of model personalization or customization [74, 25, 47]. These methods aim at embedding a new concept, e.g., pet dog, personal car, person, etc., into the output space of text-to-image models. This enables generating new images of the concept in unseen scenarios using the text prompt, e.g., my car in a field of sunflowers. To achieve this, various works fine-tune a small subset of model parameters [47, 28, 34, 89] and/or optimize text token embeddings [25, 92, 109, 2] on the few images of the new concept with different regularizations [74, 47]. More recently, encoder-based methods have been proposed that train a model on a vast dataset of concepts [81, 3, 26, 94, 48, 90, 75, 99], enabling faster customization. However, to our knowledge, no existing work allows for controlling the viewpoint in model customization. Given the ease of capturing multi-view images of a new concept, this work explores augmenting model customization with additional object viewpoint control.
@@
Figure 2: Overview. We propose a model customization method that utilizes $N$ reference images defining the 3D structure of an object $\mathcal{Y}$ (we illustrate with $N=2$ views for simplicity). We modify the diffusion model U-Net with pose-conditioned transformer blocks. Our Pose-conditioned transformer block features a FeatureNeRF module, which aggregates features from the individual viewpoints to target viewpoint $\phi$, as shown in detail in Figure 3. The rendered feature $W_{y}$ is concatenated with the target noisy feature $W_{\mathbf{x}}$ and projected to the original channel dimension. We use the diffusion U-Net itself to extract features of reference images, as shown in the top row. We only fine-tune the new parameters in linear projection layer $l$ and FeatureNerF in $F_{\text{pose}}$ blocks.
@@
View synthesis.   Novel view synthesis aims to render a scene from unseen camera poses, given multi-view images. Recently, the success of volumetric rendering-based approaches like NeRF [56] have led to numerous follow-up works with better quality [5, 6], faster speed [59, 16], and fewer training views [102, 60, 20, 86]. Recent works learn generative models with large-scale multi-view data to learn generalizable representations for novel view synthesis [50, 78, 111, 13, 95, 51, 11]. While our work draws motivation from this line of research, our goal differs - we aim to enable object viewpoint control in text-to-image personalization, rather than capturing novel views of real scenes. Concurrent to our work, ReconFusion [95] also trains a PixelNeRF [102] in the latent space of latent diffusion models for 3D reconstruction. Different from this, we learn volumetric features in the intermediate attention layers. We also focus on model customization rather than scene reconstruction. Recently, Cheng et al. [19] and Höllein et al. [33] propose adding camera pose conditioning in text-to-image diffusion models while we focus on model customization. CustomNet [104], a concurrent work, also proposes to generate custom objects in a target viewpoint in a zero-shot manner. However, it focuses primarily on generating the new object in different backgrounds, whereas our method allows any new text prompt and viewpoint combination as a condition during inference.
@@
3D editing.   Loosely related to our work, many works have been proposed for inserting and manipulating 3D objects within 2D real photographs by editing the image, using classic geometry-based approaches [39, 17, 43] or generative modeling techniques [98, 108, 55, 96, 101]. Instead of editing a single image, our work aims to “edit” the model weights of a pre-trained diffusion model. Another relevant line of work edits [29, 22] or generates [68, 88, 82, 97, 54] a 3D scene given a text prompt or image. Unlike these methods, we do not aim to edit/generate a multi-view consistent scene. Our goal is to provide additional viewpoint control when customizing text-to-image models. This enables specifying the object viewpoint while generating new backgrounds or composing multiple objects. Additionally, we show that our method achieves greater photorealism compared to a 3D editing method for this task.
@@ 3 Method
Given multi-view images of a custom object, we aim to embed it in the text-to-image diffusion model. We construct our method in order to allow the generation of new variations of the object through text prompts while providing control of the object viewpoint. Our approach involves fine-tuning the pre-trained model while conditioning it on a 3D representation of the object learned in the diffusion model’s feature space. In this section, we briefly overview the diffusion model and then explain our method in detail.
@@ 3.1 Diffusion Models
where $\mathbf{c}$ can be any modality to condition the generation process. The model is trained to predict the noise added to create the input noisy image ${\mathbf{x}}_{t}$. During inference, we gradually denoise a random Gaussian noise over a fixed number of timesteps. Various proposed sampling strategies [85, 52, 37] reduce the number of sampling steps compared to the typical $1000$ timesteps in training. In our work, we use the Stable Diffusion-XL (SDXL) [65] as the pre-trained text-to-image diffusion model. It is based on the Latent Diffusion Model (LDM) [72], which is trained in an autoencoder [45] latent space.
3.2 Customization with Object Viewpoint Control
@@
Model customization aims to condition the model on a new object, given $N$ images of the object $\mathcal{Y}=\{{\mathbf{y}}_{i}\}_{i=1}^{N}$, i.e., to model $p({\mathbf{x}}|\mathcal{Y},\mathbf{c})$ with text prompt $\mathbf{c}$. In contrast, we additionally condition the model on the object viewpoint, allowing more control in the generation process. Thus, given a set of multi-view images $\{{\mathbf{y}}_{i}\}_{i=1}^{N}$ and the corresponding camera poses $\{\pi_{i}\}_{i=1}^{N}$, our goal is to learn the conditional distribution $p({\mathbf{x}}|\{({\mathbf{y}}_{i},\pi_{i})\}_{i=1}^{N},\mathbf{c},\phi)$, where $\mathbf{c}$ is text prompt and $\phi$ is the camera pose corresponding to the target viewpoint. To achieve this, we fine-tune a pre-trained text-to-image diffusion model, which models $p({\mathbf{x}}|\mathbf{c})$, with the additional conditioning of camera pose $\phi$ given posed reference images $\{{\mathbf{y}}_{i},\pi_{i}\}_{i=1}^{N}$.
@@
Model architecture.   In Figure 2, we show the overall architecture, with an emphasis on our added pose-conditioning. Each block in the diffusion model U-Net [73] consists of a ResNet [30], denoted as $h$, followed by several transformer layers [91]. Given the output of an intermediate ResNet layer ${\mathbf{z}}$, a standard transformer layer, $F_{\text{standard}}({\mathbf{z}},\mathbf{c})$, consists of a self-attention layer, denoted as $s$, followed by cross-attention with the text prompt, denoted as $g$, and a feed-forward MLP, denoted as $f$. We modify a subset of these transformer layers to incorporate pose conditioning as we explain next.
Pose-conditioned transformer layer.   We denote the pose-
@@
conditioned transformer layer as $F_{\text{pose}}({\mathbf{z}}_{0},\{{\mathbf{z}}_{i},\pi_{i}\})$, where ${\mathbf{z}}_{0}$ is the intermediate target feature (diffusion branch in Figure 2) and $\{{\mathbf{z}}_{i}\}$ are the input features corresponding to multi-view reference images (top two rows in Figure 2). We extract spatial features $\{{\mathbf{W}}_{i}\}$ from $\{{\mathbf{z}}_{i}\}$ using components of pre-trained U-Net itself, i.e., $F_{\text{standard}}({\mathbf{z}}_{i},{\mathbf{c}})$. To condition the diffusion branch on $\phi$, we learn a radiance field, denoted as FeatureNeRF, from $\{{\mathbf{W}}_{i},\pi_{i}\}$ in a feed-forward manner [102]. The predicted FeatureNeRF is then rendered from the target viewpoint $\phi$ to obtain view-dependent feature map ${\mathbf{W}}_{y}$.
In the main diffusion branch, we extract the intermediate feature map after the self and cross-attention layers, i.e., ${\mathbf{W}}_{{\mathbf{x}}}=g(s({\mathbf{z}}_{0}),\mathbf{c})$. We concatenate ${\mathbf{W}}_{{\mathbf{x}}}$ with the rendered features ${\mathbf{W}}_{y}$ and then project it into the original feature dimension using a linear layer. Thus, the pose conditioned transformer layer, $F_{\text{pose}}({\mathbf{z}}_{0},\{{\mathbf{z}}_{i},\pi_{i}\},\mathbf{c},\phi)$ performs:
@@
Figure 3: FeatureNeRF. We predict volumetric features $\overline{\mathbf{V}}$ for each 3D point in the grid using reference features $\{{\mathbf{W}}_{i}\}$ (Eqn. 4). Given this feature, we predict the density $\sigma$ and color $rgb$ using a 2-layer MLP and use the predicted density $\sigma$ to render $\hat{\mathbf{V}}$ (which has been updated with text cross-attention $g$). The $rgb$ is only used to calculate reconstruction loss during training.
${\mathbf{W}}_{i}=F_{\text{standard}}({\mathbf{z}}_{i},$ $\mathbf{c}),\hskip 5.69054pt{\mathbf{W}}_{y}=\text{FeatureNeRF}(\{{\mathbf{W}}_{i},\pi_{i}\},\mathbf{c},\phi)$ (3)
$F_{\text{pose}}=f(l({\mathbf{W}}_{y}\oplus{\mathbf{W}}_{{\mathbf{x}}}))$
where $l$ is a learnable weight matrix, which projects the feature into the input space of feed-forward layer $f$. We initialize $l$ such that the contribution from ${\mathbf{W}}_{y}$ is zero at the start of training.
FeatureNeRF.   Here, we describe the aggregation of individual features ${\mathbf{W}}_{i}$ with poses $\pi_{i}$ into a feature map ${\mathbf{W}}_{y}$ from pose $\phi$. Given a target ray with direction $\mathbf{d}$ from target viewpoint $\phi$, we sample points $\mathbf{p}$ along the ray and project it to the image plane of each given view $\pi_{i}$. The projected coordinate is denoted as $\pi_{i}^{\mathbf{p}}$. We then sample the feature from this coordinate in ${\mathbf{W}}_{i}$, predict a feature for the 3D point $\mathbf{p}$, and aggregate the $N$ predicted features from each view with function $\psi$:
$\mathbf{V}_{i}^{\mathbf{p}}=$ $\text{MLP}(\text{Sample}({\mathbf{W}}_{i};\pi_{i}^{\mathbf{p}}),\gamma(\mathbf{d}),\gamma(\mathbf{p})),\;{i=1,...,N}$ (4)
$\overline{\mathbf{V}}^{\mathbf{p}}=\psi(\mathbf{V}_{1}^{\mathbf{p}},...,\mathbf{V}_{N}^{\mathbf{p}}),$
@@
where $\gamma$ is the frequency encoding. We use the weighted average [71] as the aggregation function $\psi$, where a linear layer predicts the weights based on $\mathbf{V}_{i}$, $\pi_{i}$, and $\phi$. For each reference view, $\mathbf{d}$ and $\mathbf{p}$ are first transformed in the view coordinate space [102]. Given the feature $\overline{\mathbf{V}}$ (superscript ${\mathbf{p}}$ is dropped for simplicity) for the 3D point, we predict the density and color using a linear layer:
$(\sigma,\mathbf{C})=\text{MLP}(\overline{\mathbf{V}}),$ (5)
@@
and also update the aggregated feature with text prompt $\mathbf{c}$ using cross-attention:
$\hat{\mathbf{V}}=\text{CrossAttn}(\overline{\mathbf{V}},\mathbf{c}).$ (6)
@@
We then render this updated feature volume using the predicted densities:
@@
where $r$ is the target ray, $\hat{\mathbf{V}}_{j}$ is the feature corresponding to the $j^{th}$ point along the ray, $\sigma_{j}$ is the predicted density of that point, $N_{f}$ is the number of sampled points along the ray between the near and far plane of the camera, and $T_{j}=\exp(-\sum_{k=1}^{j-1}\sigma_{k}\delta_{k})$ handles occlusion until that point.
We build our FeatureNeRF design based on PixelNeRF [102] but update the aggregated features with text cross-attention and use learnable weighted averaging to aggregate reference view features. Through this layer, our focus is on learning 3D features that the 2D diffusion model can use rather than learning NeRF in a feature space [100, 42].
@@
and two mask-based losses as we only wish to model the object – (1) silhouette loss [70] $\mathcal{L}_{\text{s}}$ which forces the rendered opacity to be similar to object mask, and (2) background suppression loss [7, 8] $\mathcal{L}_{\text{bg}}$ which enforces the density of all background rays to be zero.